# MARS: Reinforcing Multi-Agent Reasoning of LLMs through Self-Play in Strategic Games

### v1

@@ 4 Experiments
In this section, we present a series of extensive experiments designed to validate MARS’s ability to cultivate generalizable multi-agent skills.
@@ 4.1 Experimental setup
We use Qwen-4B (Yang et al., 2025), a state-of-the-art instruction-tuned model, as our foundation to measure the incremental multi-agent capabilities unlocked by MARS. We train two model types: specialist agents on single games (Tic-Tac-Toe, Kuhn Poker, Mini Hanabi) and a generalist agent on all three simultaneously. Our primary baseline is SPIRAL, a recent work focused on self-play in purely adversarial, zero-sum games (Liu et al., 2025). Our evaluation is structured in four stages:
@@
    Generalization to multi-agent systems: This is the primary test of this work. We integrate MARS agents into two established systems: the competitive MAD framework (Liang et al., 2023) and the cooperative AutoGen framework (Wu et al., 2024a), measuring performance on downstream math and QA benchmarks.
@@
    Pattern analysis: We conduct a qualitative analysis of the agent’s chain-of-thought to understand the emergent cognitive skills that MARS develops.
@@
Figure 4: Eval curves of Tic-Tac-Toe specialist in Tic-Tac-Toe and Mini Hanabi.
@@ 4.2 Strategic ability
Our analysis begins with strategic ability on games, with normalized game return detailed in Fig 3. The MARS framework proves highly effective: specialist agents not only master their training domains and outperform baselines, but also generalize effectively to more complex, held-out counterparts (e.g., from Tic-Tac-Toe to Connect Four). We also observe evidence of cross-category skill transfer. For example, the Tic-Tac-Toe specialist model demonstrates smooth improvement not only on Tic-Tac-Toe, but also the OOD Mini Hanabi (Fig. 4), suggesting MARS cultivates foundational skills like turn-based planning that are broadly beneficial.
Crucially, the generalist model, trained on all environments simultaneously, achieves high performance across the entire suite of games, achieving 28.7% improvement on Leduc Hold’em and 22.9% on Simple Hanabi. Its broad competence across adversarial, cooperative settings establishes it as our most robust and versatile agent overall.
Table 1: Evaluation results on downstream reasoning benchmarks within multi-agent systems. Adversarial-trained agents excel in the competitive MAD framework, while the cooperative-trained agent excels in the collaborative AutoGen framework. The generalist model performs robustly across both. Highest scores are in green; second-highest in light-green, respectively.
Setting Model Avg MATH GSM8K AQUA AIME AMC MMLU GPQA
single agent Qwen3-4B 60.74 87.60 94.60 39.80 36.70 70.00 57.10 39.39
@@
MARS (Tic.) 63.54 89.10 95.20 46.50 40.00 77.50 57.60 38.89
MARS (Kuhn.) 61.38 87.80 94.50 48.40 33.30 72.50 59.30 33.84
MARS (Hanabi) 62.05 88.10 94.70 48.00 43.30 65.00 58.90 36.36
MARS (Generalist) 62.79 89.90 94.60 52.00 33.30 75.00 59.90 34.85
MAD + Qwen3-4B 72.45 90.20 95.91 80.71 40.00 75.00 87.42 37.88
@@
MARS (Tic.) 75.01 92.20 96.06 83.07 43.33 82.50 86.76 41.12
MARS (Kuhn.) 74.54 91.60 96.21 82.68 40.00 82.50 87.39 41.41
MARS (Hanabi) 73.70 91.40 95.60 82.68 43.33 77.50 87.04 38.38
MARS (Generalist) 75.96 92.80 95.60 83.86 46.67 80.00 87.36 45.45
AutoGen + Qwen3-4B 79.14 93.40 94.69 85.04 56.67 87.50 89.21 47.47
@@
MARS (Tic.) 80.15 94.40 94.69 87.01 60.00 90.00 89.53 45.45
MARS (Kuhn.) 81.54 95.80 94.39 86.61 63.33 92.50 89.65 48.48
MARS (Hanabi) 81.54 94.40 94.54 86.22 66.67 95.00 88.98 44.95
MARS (Generalist) 82.15 95.20 94.54 86.61 66.67 92.50 89.53 50.00
@@ 4.3 Generalization to multi-agent systems
The ultimate test of MARS is whether skills honed in games transfer to practical, out-of-domain challenges. We evaluate this on a suite of demanding mathematics and general QA benchmarks, including MATH500 (Cobbe et al., 2021), GSM8K, AQUA-RAT (Ling et al., 2017), AIME24, AMC23, MMLU-STEM (Hendrycks et al., 2020), and GPQA-Diamond (Rein et al., 2024). As a preliminary step, we investigate how our multi-agent curriculum enhances foundational reasoning in a standard single-agent setting. Notably, MARS models achieve significant improvements over both the Qwen-4B baseline on a number of math benchmarks, on par with SPIRAL (Table 1).
@@
We further embed our agents into established Multi-Agent Systems (MAS) to directly measure their collaborative and competitive capabilities. In the competitive MAD debate framework, agents trained on adversarial games show a clear advantage. Conversely, in the collaborative AutoGen framework, the agents trained for cooperation—the Hanabi specialist and the generalist—excel across numerous benchmarks. In particular, the generalist model achieves a striking gain of 10% on AIME, and 12.53% on GPQA-Diamond.
@@
These results provide compelling evidence that MARS forges distinct, transferable skills for both competition and collaboration. These capabilities directly translate to improved performance in downstream multi-agent applications, with the generalist model proving to be the most robust.
4.4 Pattern analysis
@@
To understand the mechanisms behind MARS’s success, we analyze the agent’s chain-of-thought (‘<think>‘) traces. This qualitative analysis reveals the emergence of specific multi-agent skills that are cultivated in games and later transferred to general-purpose systems, as showcased in Table 2.
Table 2: Qualitative analysis of emergent cognitive patterns.
@@
First, MARS develops a Role-Aware Strategy. In Tic-Tac-Toe, the agent explicitly identifies itself as the "second-move player" and adopts a defensive strategy. This skill transfers directly to the MAD debate framework, where the agent recognizes its role as the "negative side assistant" and adopts a negative stance rather than just solving the problem. More profoundly, MARS cultivates Intent Recognition, a key component of Theory of Mind. In the cooperative Hanabi, the agent learns to interpret a teammate’s actions as communications with hidden intent (e.g., "Maybe they want me to play this card?"). This sophisticated skill transfers to the AutoGen framework, where the agent infers a collaborator’s uncertainty from subtle cues, such as a missing ‘TERMINATE‘ token, instead of treating it as a simple error.
These patterns provide qualitative evidence that MARS goes beyond improving benchmark scores; it reshapes the agent’s reasoning process, equipping it with the cognitive toolkit essential for effective multi-agent interaction.
@@ 4.5 Ablation studies
To validate our core algorithmic designs, we conduct three targeted ablation studies: (1) comparing self-play to training against a fixed opponent, and (2) sequentially removing the two key components of our advantage estimation: our "sum-then-normalize" turn-level advantage estimator and agent-specific advantage normalization. Full results are in the Appendix.
@@
Training agents against fixed, expert opponents reveals a intuitive and critical flaw: the agents overfit. As shown in Table 3, these models fail to generalize, suffering a severe performance collapse on games outside their direct training domain. This confirms that the adaptive curriculum of self-play is essential for developing robust policies.
Importance of advantage estimation design.
@@
We then ablate our two-part advantage estimation technique. First, reverting from our fine-grained turn-level advantage estimator to a coarse, trajectory-level reward signal causes a significant performance drop, especially for the Tic-Tac-Toe specialist, which is trained on a long-horizon game. Second, removing agent-specific advantage normalization also degrades performance, particularly in asymmetric games-trained model like the Tic-Tac-Toe specialist and the Kuhn Poker specialist where player experiences are distinct. Conversely, the effect is mild in the symmetric game-trained Hanabi speicialist. Together, these results confirm that both components of our advantage estimation are critical for effective learning in complex multi-agent environments.
Table 3: Average normalized game return of the fixed-opponent training ablation. The models demonstrate poor generalization, overfitting to their static opponents and performing significantly worse on out-of-distribution games. Scores below the standard MARS model are highlighted in red.
Training Games Testing Games
Model Tic. Kuhn. Mini. Con.4 Leduc. Simple.
MARS (Tic.) 75.30/32.10 42.90/-77.80 50.48 30.65/14.85 44.70/-17.48 29.75
+ fix opponent 88.00/41.95 18.60/-31.00 34.93 20.35/ 5.65 30.12/-4.65 12.22
MARS (Kuhn.) 69.85/25.50 53.70/-42.70 44.98 27.60/12.70 52.12/38.46 29.35
+ fix opponent 0.00 / 0.00 47.40/-55.30 0.00 0.00 / 0.00 -32.79/-62.38 0.00
Table 4: Average normalized game return of ablation studies on our advantage estimation technique. Both the "sum-then-normalize" approach and per-role normalization are shown to be critical for performance, as their removal leads to a notable drop in game scores. Scores below the standard MARS model are highlighted in red.
Training Games Testing Games
Model Tic. Kuhn. Mini. Con.4 Leduc. Simple.
MARS (Tic.) 75.30/32.10 42.90/-77.80 50.48 30.65/14.85 44.70/-17.48 29.75
+ w/o turn-level. 74.60/24.15 56.40/-31.90 34.80 26.75/12.30 31.40/ 4.75 19.05
+ w/o agent-specific. 82.70/31.20 35.70/-63.40 44.10 25.40/10.50 34.98/18.62 21.72
MARS (Kuhn.) 69.85/25.50 53.70/-42.70 44.98 27.60/12.70 52.12/38.46 29.35
+ w/o turn-level. 63.35/19.65 83.40/-45.40 41.65 29.60/10.85  10.04 /-11.66 22.98
+ w/o agent-specific. 69.55/24.55 45.60/-48.10 40.18 27.00/10.50  14.65 /-27.47 22.42
MARS (Hanabi) 71.90/7.35 39.30/-67.00 55.55 26.75/5.75 16.82/27.13 33.93
+ w/o turn-level. 67.55/10.60 30.30/-25.60 53.20 25.25/3.05 39.97/15.19 30.68
+ w/o agent-specific. 68.15/13.40 42.90/-65.20 52.50 32.10/5.10 26.03/29.21 32.08
@@ 5 Related work
To further extend the ability of LLMs with collaborative intelligence, LLM-based multi-agent systems have been proposed for various tasks. AutoGen (Wu et al., 2024a) and CAMEL (Li et al., 2023) designs cooperative agents for general reasoning and question answering. MetaGPT (Hong et al., 2024) and ChatDev (Qian et al., 2023) propose static multi-agent workflows with specialized roles for software development tasks. Multi-Agent Debate (Liang et al., 2023) takes a competitive approach by letting LLMs propose and criticize solutions over multiple rounds for the final answer. These works mainly focus on designing workflows for multi-agent interactions with fixed LLMs. Our work takes a complementary approach by training LLMs to enhance their multi-agent capabilities and the trained models can be integrated with these systems.
@@
Reinforcement learning (RL) has emerged as a prominent approach to enhance LLMs’ ability from instruction following to reasoning. To align LLMs with human values, RL from human feedback (RLHF) (Ouyang et al., 2022) and from AI feedback (RLAIF) (Bai et al., 2022) have been widely used in post-training to steer models towards favorable behaviors. The recent success of reasoning models (Jaech et al., 2024; Guo et al., 2025) further popularize RL from verifiable rewards (RLVR) for incentivizing long chain-of-thought reasoning capability of LLMs and have been applied in mathematical reasoning (Shao et al., 2024), coding (Wei et al., 2025), and tool-use (Jin et al., 2025). However, these works primarily consider single-turn or single agent tasks, while our approach extends RL training to multi-turn, multi-agent scenarios.
@@
The great success of achieving superhuman performance in multi-agent games (Silver et al., 2016; Vinyals et al., 2019; Berner et al., 2019) has spurred the use of self-play to train LLMs by playing with themselves. Some work (Chen et al., 2024; Wu et al., 2024b) combines game-theoretic property of self-play to overcome intransitivity of human preference in LLM alignment. Another line of work (FAIR et al., 2022; Xu et al., 2023; 2025) uses self-play to improve the strategic ability of LLM agents in specific games. The most relevant work to ours is SPAG (Cheng et al., 2024) and SPIRAL (Liu et al., 2025), which also train LLMs via self-play and show generalization to reasoning tasks. However, these work focuses on zero-sum game and show generalization in single-agent evaluation. In contrast, our work consider both cooperative and competitive games and demonstrate generalized improvement in multi-agent system.

### v2

@@ 4 Experiments
In this section, we present a series of extensive experiments designed to validate MARSHAL’s ability to obtain generalizable multi-agent skills. Additional results can be found in the Appendix.
@@ 4.1 Experimental setup
We use Qwen-4B (yang2025qwen3), a state-of-the-art instruction-tuned model, as our foundation to measure the incremental multi-agent capabilities unlocked by MARSHAL. We train two model types: specialist agents on single games (Tic-Tac-Toe, Kuhn Poker, Mini Hanabi) and a generalist agent on all three simultaneously. Our primary baseline is SPIRAL, a recent work focused on self-play in purely competitive, zero-sum games (liu2025spiral). Our evaluation is structured in four stages:
@@
    Generalization to multi-agent systems: This is the primary test of this work. We integrate MARSHAL agents into two established systems: the competitive MAD framework (liang2023encouraging) and the cooperative AutoGen framework (wu2024autogen), measuring performance on downstream math and QA benchmarks.
@@
    Pattern analysis: We conduct a qualitative analysis of the agent’s reasoning process and a quantitative analysis of failure modes to understand the successful generalization to MASs.
@@
Figure 4: Eval curves of the Tic-Tac-Toe specialist in Tic-Tac-Toe and Mini Hanabi.
@@ 4.2 Strategic ability
Our analysis begins with strategic ability on games, with normalized game return detailed in Fig 3. The MARSHAL framework proves highly effective: specialist agents not only master their training domains and outperform baselines, but also generalize effectively to more complex, held-out counterparts (e.g., from Tic-Tac-Toe to Connect Four). We also observe evidence of cross-category skill transfer. For example, the Tic-Tac-Toe specialist model demonstrates smooth improvement not only on Tic-Tac-Toe, but also the OOD Mini Hanabi (Fig. 4), suggesting MARSHAL cultivates foundational skills like turn-based planning that are broadly beneficial.
Crucially, the generalist model, trained on all environments simultaneously, achieves high performance across the entire suite of games, achieving 28.7% improvement on Leduc Hold’em and 22.9% on Simple Hanabi. Its broad competence across competitive and cooperative settings establishes it as our most robust agent overall.
Table 1: Evaluation results on downstream reasoning benchmarks within multi-agent systems. Competitive game-trained agents excel in the competitive MAD framework, while the cooperative-trained agent excels in the cooperative AutoGen framework. The generalist model performs robustly across both. Bold and underlined indicate the best and second-best scores, respectively.
Setting Model Math QA
Average MATH GSM8K AQUA AIME AMC MMLU GPQA
Qwen3-4B 60.74 87.60 94.60 39.80 36.70 70.00 57.10 39.39
@@
MARSHAL
  Tic-Tac-Toe 63.54 89.10 95.20 46.50 40.00 77.50 57.60 38.89
  Kuhn Poker 61.38 87.80 94.50 48.40 33.30 72.50 59.30 33.84
  Mini Hanabi 62.05 88.10 94.70 48.00 43.30 65.00 58.90 36.36
Single Agent Generalist 62.79 89.90 94.60 52.00 33.30 75.00 59.90 34.85
Qwen3-4B 72.45 90.20 95.91 80.71 40.00 75.00 87.42 37.88
@@
MARSHAL
  Tic-Tac-Toe 75.01 92.20 96.06 83.07 43.33 82.50 86.76 41.12
  Kuhn Poker 74.54 91.60 96.21 82.68 40.00 82.50 87.39 41.41
  Mini Hanabi 73.70 91.40 95.60 82.68 43.33 77.50 87.04 38.38
MAD (Compe- titive) Generalist 75.96 92.80 95.60 83.86 46.67 80.00 87.36 45.45
Qwen3-4B 79.14 93.40 94.69 85.04 56.67 87.50 89.21 47.47
@@
MARSHAL
  Tic-Tac-Toe 80.15 94.40 94.69 87.01 60.00 90.00 89.53 45.45
  Kuhn Poker 81.54 95.80 94.39 86.61 63.33 92.50 89.65 48.48
  Mini Hanabi 81.54 94.40 94.54 86.22 66.67 95.00 88.98 44.95
AutoGen (Coope- rative) Generalist 82.15 95.20 94.54 86.61 66.67 92.50 89.53 50.00
@@ 4.3 Generalization to multi-agent systems
The ultimate test of MARSHAL is whether skills honed in games transfer to practical, out-of-domain challenges. We evaluate this on a suite of demanding mathematics and general QA benchmarks, including MATH500 (cobbe2021training), GSM8K, AQUA-RAT (ling2017program), AIME24, AMC23, MMLU-STEM (hendrycks2020measuring), and GPQA-Diamond (rein2024gpqa). As a preliminary step, we investigate how our game tasks enhances foundational reasoning in a standard single-agent setting. Notably, MARSHAL models achieve notable improvements over both the Qwen-4B baseline on a number of math benchmarks, on par with SPIRAL (Table 1).
@@
We further embed our agents into established multi-agent systems to directly measure their cooperative and competitive capabilities. In the competitive MAD debate framework, agents trained on competitive games show a clear advantage. Notably, the generalist agent is able to achieve an average of 3.51% over the original Qwen3-4B across all benchmarks. Conversely, in the cooperative AutoGen framework, the agents trained for cooperation—the Hanabi specialist and the generalist—excel across numerous benchmarks. In particular, the generalist model achieves a striking gain of 6.57% on GPQA-Diamond in the MAD framework, and 10.00% on AIME in the AutoGen framework.
@@
These results provide compelling evidence that MARSHAL forges distinct, transferable skills for both competition and collaboration. These capabilities directly translate to improved performance in downstream multi-agent applications, with the generalist model proving to be the most robust.
4.4 Reasoning pattern and failure mode analysis
@@
To understand the mechanisms behind MARSHAL’s successful generalization, we analyze the agent’s reasoning (<think>) traces. This qualitative analysis reveals the emergence of key multi-agent skills cultivated in our mix of strategic games. We highlight two representative patterns corresponding to the competitive and cooperative training environments, as showcased in Table 2.
First, reflecting the skills honed in competitive games, MARSHAL develops a role-aware strategy. In Tic-Tac-Toe, the agent explicitly identifies itself as the "second-move player" and adopts a defensive strategy. This skill transfers directly to the MAD debate framework, where the agent recognizes its role as the "negative side assistant" and adopts a negative stance rather than just solving the problem.
More profoundly, MARSHAL cultivates intent recognition, a key component of Theory of Mind. In the cooperative Hanabi, the agent learns to interpret a teammate’s actions as communications with hidden intent (e.g., "Maybe they want me to play this card?"). This sophisticated skill transfers to the AutoGen framework, where the agent infers a collaborator’s uncertainty from subtle cues, such as a missing TERMINATE token, instead of treating it as a simple error. Together, these patterns provide qualitative evidence that MARSHAL goes beyond improving benchmark scores; it equips the agent with the cognitive toolkit essential for effective multi-agent interaction.
To further substantiate these qualitative observations with quantitative evidence, we performed a failure mode analysis on the GPQA-Diamond benchmark within the MAD framework. Adopting the taxonomy of cemri2025multi, we categorized failures into System Design Issues (e.g., formatting errors, loop repetition), Inter-Agent Misalignment (e.g., multi-agent reasoning failures like ignoring peers or task derailment), and Task Verification issues.
As shown in Fig 5 (left), while MARSHAL improves basic instruction following (reducing System Design Issues by $\sim$7%), the reduction in Inter-Agent Misalignment is significantly larger (11.5%). To understand the drivers of this strategic improvement, we further decompose the Inter-Agent Misalignment category into its sub-categories. The breakdown reveals that the performance gains are primarily driven by reductions in Task Derailment and Ignored Other Agent’s Input. This indicates that MARSHAL agent are actively listening to peers and maintaining focus on its objective, validating our hypothesis that game-theoretic self-play cultivates transferable multi-agent reasoning skills.
Table 2: Qualitative analysis of emergent reasoning patterns.
@@
Figure 5: Percentage of different failure modes in GPQA-Dimond. Through MARSHAL training, the generalist agent significantly reduces the occurrence of Inter-Agent Misalignment.
@@ 4.5 Ablation studies
To validate our core algorithmic designs, we conduct two targeted ablation studies: (1) comparing self-play to training against a fixed opponent, and (2) sequentially removing the two key algorithmic components: our turn-level advantage estimator and agent-specific advantage normalization.
@@
Training agents against fixed, expert opponents reveals a intuitive and critical flaw: the agents overfit. As shown in Table 3, these models fail to generalize, exhibiting significant performance drop on games outside their direct training domain. This is particularly evident with the Kuhn Poker specialist, which suffers a severe degradation on non-poker games—a clear case of strategic mode collapse. These results confirm that the adaptive curriculum provided by self-play is essential for the development of robust, generalizable policies.
Figure 6: Return distribution of the generalist MARSHAL agent as different player roles on Tic-Tac-Toe and Mini Hanabi.
Analysis of advantage estimation design.
@@
We then ablate our two-part advantage estimation technique. First, reverting from our fine-grained turn-level advantage estimator to a coarse, trajectory-level reward signal causes a significant performance drop, especially for the Tic-Tac-Toe specialist, which is trained on a long-horizon game. Second, removing agent-specific advantage normalization also degrades performance, particularly in competitive games-trained model like the Tic-Tac-Toe specialist and the Kuhn Poker specialist where player experiences and agent return are distinct (Fig. 6). Conversely, the effect is mild in the Hanabi specialist, which is trained on a cooperative game with similar return distribution between players (Fig. 6). Together, these results confirm that both components of our advantage estimation are critical for effective learning in complex multi-turn, multi-agent environments.
Table 3: Generalization comparison between MARSHAL (self-play) and its fixed-opponent variant. The latter exhibits significant overfitting to static environments and opponents. Values denote average normalized game returns; for competitive games, entries indicate first-move / second-move returns. Underlined scores indicate performance degradation compared to the standard MARSHAL model.
Model Training Games Testing Games
Tic-Tac-Toe Kuhn Poker
Mini
Hanabi
Connect Four
Leduc
Hold’em
Simple
Hanabi
MARSHAL (Tic-Tac-Toe) 75.30 / 32.10 74.15 / 3.42 50.48 30.65 / 14.85 58.36 / 27.65 29.75
w/ fixed opponent 88.00 / 41.95 63.15 / 28.84 34.93 20.35 / 5.65 47.38 / 35.55 12.22
MARSHAL (Kuhn Poker) 69.85 / 25.50 79.04 / 22.49 44.98 27.60 / 12.70 63.94 / 62.10 29.35
w/ fixed opponent 0.00 / 0.00 76.19 / 15.64 0.00 0.00 / 0.00 0.00 / 0.00 0.00
Table 4: Ablation results for algorithmic design. Both our turn-level advantage estimator and agent-specific advantage normalization prove essential for performance. Notation follows Table 3.
Model Training Games Testing Games
Tic-Tac-Toe Kuhn Poker
Mini
Hanabi
Connect Four
Leduc
Hold’em
Simple
Hanabi
MARSHAL (Tic-Tac-Toe) 75.30 / 32.10 74.15 / 3.42 50.48 30.65 / 14.85 58.36 / 27.65 29.75
w/o Turn-Level. 74.60 / 24.15 80.26 / 28.35 34.80 26.75 / 12.30 48.34 / 41.34 19.05
w/o Agent-Specific. 82.70 / 31.20 70.89 / 11.24 44.10 25.40 / 10.50 51.04 / 49.88 21.72
MARSHAL (Kuhn Poker) 69.85 / 25.50 79.04 / 22.49 44.98 27.60 / 12.70 63.94 / 62.10 29.35
w/o Turn-Level. 63.35 / 19.65 92.49 / 21.02 41.65 29.60 / 10.85 32.26 / 31.23 22.98
w/o Agent-Specific. 69.55 / 24.55 75.37 / 19.55 40.18 27.00 / 10.50 35.73 / 21.50 22.42
MARSHAL (Hanabi) 71.90 / 7.35 72.52 / 9.29 55.55 26.75 / 5.75 37.36 / 55.12 33.93
w/o Turn-Level. 67.55 / 10.60 68.45 / 31.78 53.20 25.25 / 3.05 54.79 / 47.77 30.68
w/o Agent-Specific. 68.15 / 13.40 74.15 / 10.27 52.50 32.10 / 5.10 44.30 / 56.41 32.08
@@ 5 Related work
To further extend the ability of LLMs with collaborative intelligence, LLM-based multi-agent systems have been proposed for various tasks. AutoGen (wu2024autogen) and CAMEL (li2023camel) designs cooperative agents for general reasoning and question answering. MetaGPT (hong2024metagpt) and ChatDev (qian2023chatdev) propose static multi-agent workflows with specialized roles for software development tasks. Multi-Agent Debate (liang2023encouraging) takes a competitive approach by letting LLMs propose and criticize solutions over multiple rounds for the final answer. These works mainly focus on designing workflows for multi-agent interactions with fixed LLMs. Other work have proposed to generate workflows automatically (zhang2024aflow). Our work takes a complementary approach by training LLMs to enhance their multi-agent capabilities and build stronger foundation models that can be integrated with these systems.
@@
Reinforcement learning (RL) has emerged as a prominent approach to enhance LLMs’ ability from instruction following to reasoning. To align LLMs with human values, RL from human feedback (RLHF) (ouyang2022training) and from AI feedback (RLAIF) (bai2022constitutional) have been widely used in post-training to steer models towards favorable behaviors. The recent success of reasoning models (jaech2024openai; guo2025deepseek) further popularize RL from verifiable rewards (RLVR) for incentivizing long Chain-of-Thought reasoning capability of LLMs and have been applied in mathematical reasoning (shao2024deepseekmath), coding (wei2025swe), and tool-using (jin2025search). However, these works primarily consider single-turn or single-agent tasks, while our approach extends RL training to multi-turn, multi-agent scenarios. Recently, MT-GRPO (zeng2025reinforcing) also address the turn-level credit assignment issue for multi-turn RL. However, they propose to normalize the turn-level rewards turn-by-turn, which can introduce variance in strategic games with varying number of turns. We propose a "sum-then-normalize" approach that takes advantage of all turns in the normalization, increasing robustness.
@@
The great success of achieving superhuman performance in multi-agent games (silver2016mastering; vinyals2019grandmaster; berner2019dota) has spurred the use of self-play to train LLMs by playing with themselves (zhang2024survey). Some work (chen2024self; wu2024self) combines game-theoretic property of self-play to overcome intransitivity of human preference in LLM alignment. Another line of work (meta2022human; xu2023language; xu2025learning) uses self-play to improve the strategic ability of LLM agents in specific games. The most relevant work to ours is SPAG (cheng2024self) and SPIRAL (liu2025spiral), which also train LLMs via self-play to show generalization to reasoning tasks. However, these work focuses on zero-sum game and show generalization in single-agent evaluations. In contrast, our work consider both cooperative and competitive games and demonstrate generalized improvement in multi-agent system. As a concurrent work to ours, SPIRAL also address the heterogeneity of player role in advantage estimation and introduces a Role-Conditioned Advantage Estimation (RAE), analogous to our agent-specific normalization. We advance this finding by revealing that role separation is only critical in games with distinct return distributions between players.